# Going beyond the creative limits of next-token prediction

### v2

@@ Abstract
We design a suite of minimal algorithmic tasks that are a loose abstraction of open-ended real-world tasks. This allows us to cleanly and controllably quantify the creative limits of the present-day language model. Much like real-world tasks that require a creative, far-sighted leap of thought, our tasks require an implicit, open-ended stochastic planning step that either (a) discovers new connections in an abstract knowledge graph (like in wordplay, drawing analogies, or research) or (b) constructs new patterns (like in designing math problems or new proteins). In these tasks, we empirically and conceptually argue how next-token learning is myopic and memorizes excessively; multi-token approaches, namely teacherless training and diffusion models, comparatively excel in producing diverse and original output. Secondly, to elicit randomness without hurting coherence, we find that injecting noise at the input layer (dubbed as seed-conditioning) works surprisingly as well as (and in some conditions, better than) temperature sampling from the output layer. Thus, our work offers a principled, minimal test-bed for analyzing open-ended creative skills, and offers new arguments for going beyond next-token learning and temperature sampling. We make part of the code available under https://github.com/chenwu98/algorithmic-creativity
@@ 1 Introduction
In search of more definitive conclusions, we approach from a different angle: we study minimal and controllable tasks that are loose abstractions of real-world tasks and yet allow one to rigorously quantify originality and diversity. This follows along the lines of recent works that have studied the diversity of models in graph path-finding (Khona et al., 2024) and generating challenging CFGs (Allen-Zhu & Li, 2023b). Broadly, we refer to such tasks as open-ended algorithmic tasks. Our aim is to design tasks more minimal than these prior tasks, and crucially, tease apart distinct computational skills required for creativity. This will allow us to systematically investigate issues in the current paradigm of model training and propose alternatives.
As our first main contribution, we draw inspiration from cognitive science literature (Boden, 2003) (see also Franceschelli & Musolesi (2023)) to design algorithmic tasks isolating two distinct types of creative leaps of thought. The first class of tasks involves combinational creativity: drawing novel connections in knowledge, like in research, wordplay or drawing analogies (see Fig 1 for task description). The second class of tasks involves exploratory creativity: constructing fresh patterns subject to certain rules, like in designing problems and suspense (see Fig 2). In these tasks, we can precisely evaluate models for the fraction of generations that are coherent, unique and original (not present in training set). We term this metric “algorithmic creativity” to denote that it is solely evaluates the computational aspects of creativity.
@@
Within this framework, we articulate two creative limits of the current language modeling paradigm. First, we empirically find that next-token learning achieves lower algorithmic creativity (and higher memorization) compared to multi-token approaches, namely, teacherless training (Bachmann & Nagarajan, 2024; Monea et al., 2023; Tschannen et al., 2023) and diffusion models (Hoogeboom et al., 2021; Austin et al., 2021; Lou et al., 2023) (see Fig 3 and Fig 4). Our argument is that in all our tasks, inferring the latent leap of thought requires observing global higher-order patterns rather than local next-token patterns in the sequence.
Next, we turn to how we elicit randomness from a Transformer. While the de facto approach is to elicit randomness at the output — temperature sampling — we contrast this with injecting randomness into the model. Concretely, we study seed-conditioning, where we train with a random prefix string per example. Surprisingly, we find that seed-conditioning induces non-trivial algorithmic creativity even with deterministic, greedy decoding. In fact, seed-conditioning is competitive with and in some cases, more creative than temperature sampling. Intuitively, maximizing diversity at the output-token-level is computationally burdensome: it requires simultaneously processing a diverse set of leaps of thoughts to compute a marginalized token distribution. It is easier to first sample a single latent leap of thought, and then compute the token conditioned on that one leap. We conjecture that seed-conditioning enables a latent-leap-conditioned token generation.
@@
    We find that multi-token prediction through teacherless training or diffusion, results in significantly increased algorithmic creativity and reduced memorization in our tasks compared to next-token prediction.
@@
    Our argument provides new support for multi-token prediction, going beyond (B&N’24). We show a gap in creativity in an open-ended task (rather than correctness in a deterministic one), in much simpler 2-token-lookahead tasks, and in tasks where no token permutation is friendly to next-token-learning.
@@
    We find that input-randomization via seed-conditioning achieves algorithmic creativity competitive with conventional output-randomization via temperature sampling.
@@ 2 Open-ended algorithmic tasks & two types of creativity
Open-ended tasks that do not require a leap of thought. One simple open-ended task that may come to mind is generating uniformly-random known entities, like celebrity names (Zhang et al., 2024b). However, there is no opportunity to create a novel string here. A more interesting example may be generating grammatically coherent PCFG strings following a subject verb object format e.g., the cat chased a rat (Hopkins et al., 2023). While novel strings become possible here, no sophisticated leaps of thought are involved; each token can be generated on the fly, satisfying a local next-token constraint to be coherent.
@@ 2.1 The fundamental types of creativity (Boden, 2003)
In the upcoming sections, we will attempt to capture some core computations of rudimentary instances within the two classes of creative skills above. We emphasize that by no means does our minimal algorithmic setup intend to capture the human values that go into these endeavors; nor do they capture the rich array of creative acts that Boden (2003) discusses within these categories. (See limitations in §6)
@@ 2.2 The basic setting and notations
In all our tasks, we assume the standard generative model setting: the model must learn an underlying distribution $\mathcal{D}$ through a training set $S$ of $m$ independent samples $\bm{s}_{i}\sim\mathcal{D}$. The distribution is over a space $\mathbb{V}^{L}$ of $L$-length strings. The tasks are open-ended in that there is no one correct answer at test-time. The goal is to produce a random string from $\mathcal{D}$, much like responding to the query Design a high-school word problem.
@@ 3 Training and Inference
For our next-token-trained (NTP) models, we use the standard teacher-forcing objective used in supervised finetuning. Given prompt $\bm{p}$ and ground truth sequence $\bm{s}$, the model is trained to predict the $i$’th token $s_{i}$, given as input the prompt and all ground truth tokens up until that point, $(\bm{p},\bm{s}_{<i})$. We write the objective more explicitly in §A Eq 2. For the multi-token Transformer models, we use teacherless training (Monea et al., 2023; Bachmann & Nagarajan, 2024; Tschannen et al., 2023), where the model is trained to predict $s_{i}$ simultaneously for all $i$, only given the prompt $\bm{p}$ (and some dummy tokens in place of the $\bm{s}$ that was once given as input). Since the exact details of this is irrelevant to our discussion, we describe this in Eq 2. To train our models, we use a hybrid of this objective and the next-token objective.
@@
Rather than sequentially predicting each token conditioned on previously generated tokens, discrete diffusion models (Hoogeboom et al., 2021; Austin et al., 2021) iteratively add noise to all tokens and then learn to denoise them in reverse. This strategy allows the model to capture global dependencies among tokens during training, making it an example of a multi-token objective. In our experiments, we used the score entropy discrete diffusion model (SEDD, Lou et al., 2023), which starts generation from a sequence of fully masked tokens. Over multiple steps, the model simultaneously predicts and unmasks these tokens, progressively refining the entire sequence.
@@ 4.1 Observations
Figure 4: Multi-token diffusion training improves algorithmic creativity (top; Eq 1) on our four open-ended tasks, and reduces memorization on discovery tasks but not construction tasks (bottom).
In Fig 4, we report performance for the diffusion model against next-token & teacherless training of similar-sized Transformers. We see that diffusion models are consistently better than next-token training, achieving up to 5x higher algorithmic creativity. However, the gains are much smaller or absent with teacherless training. This echoes prior discussions suggesting that the teacherless objective is a hard objective to optimize for smaller Transformers (B&N’24); other multi-token approaches (Gloeckle et al., 2024) are even known to hurt small models of the order of 300M.
Multi-token prediction reduces memorization significantly. Algorithmic creativity may suffer either because the model outputs incoherent garbage, or because it repeats the same original output, or because it simply parrots out the training data. In almost all settings, it is the last reason that dominates: across the board (in Fig 3, Fig 4 bottom), next-token prediction is significantly prone to memorizing the data, while multi-token methods are highly resistant. As foreshadowed in §2.6, we hypothesize that this is because NTP memorizes the earlier training tokens without a global plan, having fit the later tokens via local coherence rules (because of Clever Hans cheats à la B&N’24). Note that an exception to this is the smaller models (especially for diffusion) in our construction tasks, where memorization increases under the multi-token objectives; but this increase is mild and crucially, does not hurt algorithmic creativity.
@@
Figure 6: Seed-conditioning improves algorithmic creativity of the GPT-2 (86M) model: All models are trained with NTP. The color denotes temperature, while the X-axis, the prefix. Remarkably, seed-conditioning achieves high creativity even with greedy decoding. However, note that for the discovery datasets, top-K sampling must be enabled to see a positive gap between seed- and null- settings.
@@
Next, in comparison to temperature sampling, in our settings, seed-conditioning results in the highest algorithmic creativity in both the small models (Fig 6) and the larger models (Fig 5). In fact, in our larger models, the null and pause token prefix with temperature sampling (Fig 18) exhibit almost no algorithmic creativity (they are mode collapsed, see Fig 1920). Besides, for any fixed temperature, prefixing a seed only improves performance over a null prefix (Fig 618).
In §H.2, we find that the improved creativity from seed-conditioning comes from improved diversity, rather than from reduced memorization (unlike with multi-token training). In some of our settings, this positive gap between seed-conditioning and temperature sampling is attributable to the fact that top-$K$ sampling is enabled by default as is standard (specifically, for GPT-2 (86M) on Sibling Discovery and Triangle Discovery where the vocabulary size is larger than $K$). Intuitively, top-$K$ sampling restricts the degree of freedom in output randomness, allowing the seed to introduce new degrees of freedom. However, in all other settings, this explanation does not hold. Regardless all these behaviors of seed-conditioning require further study as they are non-trivial, especially given that the seeds are arbitrarily chosen in relationship to the output. As a starting point, we provide some preliminary hypotheses for when and why seed-conditioning may help Transformers in §5.1. Note that we do not see gains of seed-conditioning when it comes to diffusion training (§ F.3).
@@ 5 Discussion
5.1 Intuition about seed-conditioning
Seed-conditioning produces non-trivial algorithmic creativity even with greedy decoding. Perhaps, one could view the seed prefixes as a simpler alternative to varying the wordings of a prompt Li et al. (2023); Lau et al. (2024); Naik et al. (2024) or tuning a soft-prompt (Wang et al., 2024c), both of which are known to induce diversity. However, it is unexpected that the seeds provide any meaningful randomness at all, given they are arbitrarily picked with no semantic relationship to the training point. We do not understand this presently.
Another pertinent question is to why seed-conditioning works better than temperature sampling in the settings it does. To this, we offer three potential explanations. The simplest explanation is that in our GPT-2 (86M) settings, top-K sampling is enabled by default. This restricts the degree of freedom in the support induced by the output-layer randomness; thus, providing different seeds can help vary this support and produce more varying samples. However, this does not explain the gap we see in the other datasets where the output vocabulary is smaller than $K$ ($K=50$), or in Gemma v1 (2B) where we do not use top-K sampling.
@@ 5.1 Intuition about seed-conditioning
Our argument is limited to learning open-ended tasks in a supervised manner. While we do not comment on how well other approaches like RL (DeepSeek-AI, 2025), chain-of-thought (CoT; Wei et al., 2022), and scaling test-time compute (OpenAI, 2024) would fare, we remark that these methods are designed to enhance the quality of a single example. It is unclear how to design them to maximize originality against a training set, and diversity over multiple responses. Furthermore, there is a profound question as to whether merely spelling out a model’s thought in token space can be an efficient way to search and maximize diversity. This may require enumerating all possible candidates by trial and error, an impossible feat when the search space is large.
@@ 6.2 Our approach to creativity
    As stated earlier, we study abstract tasks that are inspired by the computations involved in creative tasks. Our study is not intended to capture the subjective, social, cultural and personal values integral to many creative tasks.
@@ C.1 Datasets inspired by combinational creativity
Dataset 1: Sibling Discovery.
@@
Dataset 2: Triangle Discovery
@@ C.2 Datasets inspired by exploratory creativity
Dataset 3: Circle Construction.
@@
Dataset 4: Line Construction
@@ Appendix D Further experimental details
Details for GPT-2 (86M) model.  We use GPT-2 (small) with 86M non-embedding parameters when we are comparing Transformers with diffusion models. We train these models with a learning rate of $10^{-4}$ and a batch size of $64$, to convergence in terms of the algorithmic creativity. In temperature sampling, we use nucleus sampling with a parameter of $0.95$, although we observed this does not affect the results qualitatively. We used top-k sampling with a parameter of $50$. We also explore temperatures above $1.0$ for these models since temperatures low than $1.0$ only did worse. We provide sensitivity analysis of learning rate in §F.

### v3

@@ Abstract
We design a suite of minimal algorithmic tasks that are a loose abstraction of open-ended real-world tasks. This allows us to cleanly and controllably quantify the creative limits of the present-day language model. Much like real-world tasks that require a creative, far-sighted leap of thought, our tasks require an implicit, open-ended stochastic planning step that either (a) discovers new connections in an abstract knowledge graph (like in wordplay, drawing analogies, or research) or (b) constructs new patterns (like in designing math problems or new proteins). In these tasks, we empirically and conceptually argue how next-token learning is myopic; multi-token approaches, namely teacherless training and diffusion models, comparatively excel in producing diverse and original output. Secondly, to elicit randomness without hurting coherence, we find that injecting noise at the input layer (dubbed seed-conditioning) works surprisingly as well as (and in some conditions, better than) temperature sampling from the output layer. Thus, our work offers a principled, minimal test-bed for analyzing open-ended creative skills, and offers new arguments for going beyond next-token learning and temperature sampling. We make part of the code available under https://github.com/chenwu98/algorithmic-creativity
@@ 1 Introduction
In search of more definitive conclusions, we approach from a different angle: we study minimal and controllable tasks that are loose abstractions of real-world tasks and yet allow one to rigorously quantify originality and diversity. This follows along the lines of recent works that have studied the diversity of models in graph path-finding (Khona et al., 2024) and generating challenging context-free grammars (CFGs) (Allen-Zhu & Li, 2023b). Broadly, we refer to such tasks as open-ended algorithmic tasks. Our aim is to design tasks more minimal than these prior tasks, and crucially, tease apart distinct computational skills required for creativity. This will allow us to systematically investigate issues in the current paradigm of model training and propose alternatives.
As our first main contribution, we draw inspiration from cognitive science literature (Boden, 2003) (see also Franceschelli & Musolesi (2023)) to design algorithmic tasks isolating two distinct types of creative leaps of thought. The first class of tasks involves combinational creativity: drawing novel connections in knowledge, like in research, wordplay or drawing analogies (see Fig 1 for task description). The second class of tasks involves exploratory creativity: constructing fresh patterns subject to certain rules, like in designing problems and suspense (see Fig 2). In these tasks, we can precisely evaluate models for the fraction of generations that are coherent, unique and original (not present in training set). We term this metric “algorithmic creativity” to denote that it solely evaluates the computational aspects of creativity.
@@
Within this framework, we articulate two creative limits of the current language modeling paradigm. First, we empirically find that next-token learning achieves lower algorithmic creativity compared to some multi-token approach, namely, either teacherless training (Bachmann & Nagarajan, 2024; Monea et al., 2023; Tschannen et al., 2023) or diffusion (Hoogeboom et al., 2021; Austin et al., 2021; Lou et al., 2023) (see Fig 3 and Fig 4). Our argument is that in all our tasks, inferring the latent leap of thought requires observing global higher-order patterns rather than local next-token patterns in the sequence.
Next, we turn to how we elicit randomness from a Transformer. While the de facto approach is to elicit randomness at the output — temperature sampling — we contrast this with injecting randomness into the model. Concretely, we study seed-conditioning, where we train with a random prefix string per example. Surprisingly, we find that seed-conditioning induces non-trivial algorithmic creativity even with deterministic, greedy decoding. In fact, seed-conditioning is comparable with and in some cases, more creative than temperature sampling. Intuitively, maximizing diversity at the output-token-level is computationally burdensome: it requires simultaneously processing a diverse set of leaps of thoughts to compute a marginalized token distribution. It is easier to first sample a single latent leap of thought, and then compute the token conditioned on that one leap. We conjecture that seed-conditioning enables a latent-leap-conditioned token generation.
@@
    We find that multi-token prediction through teacherless training or diffusion, results in significantly increased algorithmic creativity in our tasks compared to next-token prediction.
@@
    Our argument provides new support for multi-token prediction, going beyond B&N’24. We show a gap in creativity in an open-ended task (rather than correctness in a deterministic one), in much simpler 2-token-lookahead tasks, and in tasks where no token permutation is friendly to next-token-learning.
@@
    We find that input-randomization via seed-conditioning achieves algorithmic creativity comparable with conventional output-randomization via temperature sampling.
@@ 2 Open-ended algorithmic tasks & two types of creativity
Open-ended tasks that do not require a leap of thought. One simple open-ended task that may come to mind is generating uniformly-random known entities, like celebrity names (Zhang et al., 2024b). However, there is no opportunity to create a novel string here. A more interesting example may be generating grammatically coherent probabilistic CFG strings following a subject verb object format e.g., the cat chased a rat (Hopkins et al., 2023). While novel strings become possible now, no sophisticated leaps of thought are involved; each token can be generated on the fly, satisfying a local next-token constraint to be coherent.
@@ 2.1 The fundamental types of creativity (Boden, 2003)
In the upcoming sections, we will attempt to capture some core computations of rudimentary instances within the two classes of creative skills above. We emphasize that by no means does our minimal algorithmic setup intend to capture the human values or the subjective aspects of these endeavors (Schmidhuber, 2009); nor do they capture the rich array of creative acts that Boden (2003) discusses within these categories. (See limitations in §6).
@@ 2.2 The basic setting and notations
In all our tasks, we assume the standard generative model setting: the model must learn an underlying distribution $\mathcal{D}$ through a training set $S$ of $m$ independent samples $\bm{s}_{i}\sim\mathcal{D}$. The distribution is over a space $\mathbb{V}^{L}$ of $L$-length strings composed of tokens from the vocabulary $\mathbb{V}$. The tasks are open-ended in that there is no one correct answer at test-time. The goal is to produce a random string from $\mathcal{D}$, much like responding to the query Design a high-school word problem. We note that this setup resembles how models are trained in pretraining or in applications like drug discovery or protein modeling.
@@ 3 Training and Inference
For our next-token-trained (NTP) models, we use the standard teacher-forcing objective used in supervised finetuning. Given prompt $\bm{p}$ and ground truth sequence $\bm{s}$, the model is trained to predict the $i$’th token $s_{i}$, given as input the prompt and all ground truth tokens up until that point, $(\bm{p},\bm{s}_{<i})$. We write the objective more explicitly in §A Eq 2. For the multi-token Transformer models, we use teacherless training (Monea et al., 2023; Bachmann & Nagarajan, 2024; Tschannen et al., 2023), where the model is trained to predict $s_{i}$ simultaneously for all $i$, only given the prompt $\bm{p}$ (and some dummy tokens masking the $\bm{s}$ that was once given as input). Since the exact details of this is irrelevant to our discussion, we describe this in Eq 2. To train our models, we use a hybrid of this objective and the next-token objective.
@@
In our experiments, we use the score entropy discrete diffusion model (SEDD, Lou et al., 2023). Loosely, we can think of the objective here as a hybrid of the teacherless objective, where the input is fully masked or corrupted (and the model must unmask or uncorrupt all of it), and various easier objectives with partial maskings or corruptions of the input. During inference, the model starts with a fully masked or corrupted sequence of tokens, and iteratively corrects subsets of them.
@@ 4.1 Observations
Figure 4: Multi-token diffusion training improves algorithmic creativity (top; Eq 1) on three of our four open-ended tasks on GPT-2 (86M) and similarly-sized diffusion model, SEDD (90M). We report the best performance after tuning the sampling hyperparameters – temperature from $\{0,0.5,1,2\}$, with or without top-$K$ where $K=50$.
For the smaller models, in Fig 4, we find that the diffusion model improves over next-token training of similar-sized Transformers on all tasks, except on Sibling Discovery where it is mildly worse. With teacherless training, the gains are absent; however we did find that the creativity of the top-$K$ samples is generally improved (Fig 24), indicating a less peaky distribution. Regardless, this echoes prior discussions suggesting that for smaller Transformers, teacherless training is hard to optimize (B&N’24) and other multi-token objectives even hurt (Gloeckle et al., 2024) smaller models of the order of 300M.
Multi-token prediction reduces memorization for the larger model. Algorithmic creativity may suffer either because the model outputs incoherent garbage, or because it repeats the same original output, or because it simply parrots out the training data. For our larger model, it is the last reason that dominates: across the board (in Fig 3), next-token prediction is significantly prone to memorizing the data, while the multi-token method is highly resistant. As foreshadowed in §2.6, we hypothesize that this is because NTP memorizes the earlier training tokens without a global plan, having fit the later tokens via local coherence rules (because of Clever Hans cheats à la B&N’24). Note that the smaller models have much lower absolute memorization values; the multi-token objective preserves or mildly worsens this, in a way that does not always hurt creativity proportionally. We did find that amongst the top-$K$ samples, teacherless training has reduced memorization for some tasks (see Fig 25).
@@
Figure 6: Seed-conditioning achieves high algorithmic creativity in the GPT-2 (86M) model: All models are trained with NTP. The color denotes temperature, while the X-axis, the prefix. Remarkably, seed-conditioning achieves high creativity even with greedy decoding, and improves creativity on all except Sibling Discovery.
@@
Next, in comparison to temperature sampling, in all our settings, seed-conditioning results in the algorithmic creativity comparable to temperature sampling. In fact, for the larger models (Fig 5), seed-conditioning generally outperforms the baseline “null” conditioning (which appears mode-collapsed, see Fig 181920). For the smaller model as well (Fig 6), we find gains in all tasks except for Sibling Discovery. Besides, for any fixed temperature, prefixing a seed only improves performance over a null prefix (Fig 618).
In §H.2, we find that the improved creativity from seed-conditioning comes from improved diversity, rather than from reduced memorization (unlike with multi-token training). All these behaviors of seed-conditioning require further study as they are non-trivial, especially given that the seeds are naively, arbitrarily chosen in relationship to the output. As a starting point, we provide some preliminary hypotheses for when and why seed-conditioning may help Transformers in §5.1. Note that we do not see gains of seed-conditioning when it comes to diffusion training (§ F.3).
@@ 5 Discussion
5.1 Intuition for seed-conditioning
Seed-conditioning produces non-trivial algorithmic creativity even with greedy decoding. Perhaps, one could view the seed prefixes as a simpler alternative to varying the wordings of a prompt (Li et al., 2023; Lau et al., 2024; Naik et al., 2024) or tuning a soft-prompt (Wang et al., 2024c), both of which are known to induce diversity. However, it is unexpected that the seeds provide any meaningful randomness at all, given they are arbitrarily picked with no semantic relationship to the training point. We do not understand this presently.
Another pertinent question is to why seed-conditioning works better than temperature sampling in the settings it does. To this, we offer three potential explanations. The simplest explanation is that in the initial version of our paper, top-K sampling was enabled by default with $K=50$ for the GPT-2 (86M) settings. This would have restricted the degree of freedom in the support induced by the output-layer randomness; thus, providing different seeds can help vary this support and produce more varying samples. However, even upon removing top-K sampling, we see gains in our exploratory creativity tasks for GPT-2 (86M) (after all their vocabulary is smaller than $K=50$), and in all our Gemma v1 (2B) results. This is unexplained.
@@ 5.1 Intuition for seed-conditioning
Our argument is limited to learning open-ended tasks in a supervised manner which directly corresponds to how we perform pretraining or apply models for protein or drug discovery. While we do not comment on how well other powerful approaches like RL (DeepSeek-AI, 2025), chain-of-thought (CoT; Wei et al., 2022), and scaling test-time compute (OpenAI, 2024) would address the limitations we bring up, we make a few remarks. First, in cases where post-training only elicits pre-existing skills in the base model (Muennighoff et al., 2025), our limitations and suggestions about a next-token-trained base model remain relevant. Next, regardless of the benefits one may get from pre-training, our arguments imply that pre-training squanders its rich supervision, when it could be vastly compute- and data-efficient. Finally, there is a profound question as to whether exploration through sparse rewards can learn the creative process during training-time or whether articulation in the token space (with a thinking model) can efficiently maximize diversity during test-time (which may require brute-force enumeration of all candidates).
@@ 6.1 Limitations of our experimental conclusions
  7. 7.
    There are many ways in which our observations have not been fully characterized or understood even in our minimal settings. For example, we do not thoroughly report on the effect of model size, model family, or pretraining; we do not analyze when and why seed-conditioning helps.
@@ 6.2 Our approach to creativity
    As stated earlier, we study abstract tasks that are inspired by the computations involved in creative tasks. Our study is not intended to capture the subjective aspects like interestingness (Schmidhuber, 2009) and other social, cultural and personal values integral to many creative tasks.
@@ C.1 Datasets inspired by combinational creativity
Dataset 1: Sibling Discovery (Fig 9(a)).
@@
Dataset 2: Triangle Discovery (Fig 9(b)).
@@ C.2 Datasets inspired by exploratory creativity
Dataset 3: Circle Construction (Fig 10(a)).
@@
Dataset 4: Line Construction (Fig 10(a)).
@@ Appendix D Further experimental details
Details for GPT-2 (86M) model. In Table 2, we provide the hyperparameter details for each of our datasets.  We use GPT-2 (small) with 86M non-embedding parameters when we are comparing Transformers with diffusion models. We train these models with a learning rate of $10^{-4}$ and a batch size of $64$, to convergence in terms of the algorithmic creativity. In temperature sampling, we use nucleus sampling with a parameter of $0.95$, although we observed this does not affect the results qualitatively. We used top-k sampling with a parameter of $50$. We also explore temperatures above $1.0$ for these models since temperatures low than $1.0$ only did worse. We provide sensitivity analysis of learning rate in §F.
Table 2: Hyperparameter details for GPT-2 (86M) and SEDD (90M) model.
Max. Learning Rate $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
Model Seq. Len. $128$ $128$ $128$ $128$
Max training steps (GPT-2 (86M)) $20k$ $80k$ $40k$ $40k$
Max training steps (SEDD (90M)) $80k$ $80k$ $80k$ $80k$
Training size $50k$ $15k$ $10k$ $10k$
Weight given to multi-token obj. $0.5$ $0.33$ $0.75$ $0.75$